EXPLAIN WITH THE HELP OF GRAPH, A monopolist has a cost


function 200q + 15Q2


and faces a demand function given by P= 1200 – 10Q. Calculate total


revenue, marginal revenue, output and price that maximize profit-maximizing? What is its


maximal profit?

1
Expert's answer
2021-12-15T16:20:48-0500

Total revenue is:

"TR = P\u00d7Q = 1200Q - 10Q^2."

Marginal revenue is:

"MR = TR'(Q) = 1200 - 20Q."

At MR = MC output and price maximize profit, so:

"MC = TC'(Q) = 200 + 30Q,"

1200 - 20Q = 200 + 30Q,

50Q = 1000,

Q = 20 units,

P = 1200 - 10×20 = 1000.

Total profit is:

"TP = TR - TC = 1000\u00d720 - (200\u00d720 + 12\u00d720^2) = 11200."